OpenStudy (anonymous):

find the area of the region containing points (x,y) satisfying sin(x+y)>=0 and x^2+y^2<=100.

OpenStudy (nikvist):

\[S_{area}=\frac{S_{circle}}{2}=\frac{10^2\pi}{2}=50\pi\]

OpenStudy (anonymous):

Dont think its that simple^

OpenStudy (anonymous):

the inequality with the sin gives a series of strips

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin%28x%2By%29+%3E%3D+0%2C++x%5E2+%2By%5E2%3C%3D100 much more complicated region , I am not even sure if its possible

OpenStudy (nikvist):

Sum of all regions is semicircle